# The Identification Power of

### v1

@@ Appendix C Proofs
We will show that $F\in\widetilde{\mathcal{F}}^{\ast}$ and $\mathbb{E}_{F}[\psi(Y_{1},Y_{0})]=\theta$. For the former, we will specifically show that (i) $F$ satisfies Assumptions 2.12.5 and 5.15.2 (with $F^{\ast}$ replaced by $F$) and that (ii) $F_{YDGX}=F_{YDGX}^{\ast}$.
We begin with verifying condition (i). By the same argument as in the proof of Theorem 3.2, we can show that $F$ satisfies Assumptions 2.12.5, or equivalently $F\in\mathcal{F}^{\ast}$. Applying Lemma C.1 (i) to (56), we have
$F_{Y_{1}Y_{0}|SX}(\cdot,\cdot|S,X)=C\left(F_{Y_{1}|SX}^{\ast}(\cdot|S,X),F_{Y_{0}|SX}^{\ast}(\cdot|S,X)\middle|S,X\right).$
Then, by definition of $\widetilde{\mathcal{C}}$ and the equation $F_{Y_{d}SX}=F_{Y_{d}SX}^{\ast}$, which follows from Lemma C.1 and (55), we obtain $F_{Y_{1}Y_{0}SX}\in\widetilde{\mathcal{F}}_{Y_{1}Y_{0}SX}^{\dagger}$. Because $F_{Y_{d}SX}=F_{Y_{d}SX}^{\ast}$, conditions (15) and (15) are equivalent to Assumptions 5.1 and 5.2, respectively. Hence $F$ satisfies Assumptions 5.1 and 5.2.
To sum up, we have shown that $F$ satisfies condition (i) (i.e., $F$ satisfies Assumptions 2.12.5 and 5.15.2 with $F^{\ast}$ replaced by $F$). Condition (ii) (i.e., $F_{YDGX}=F_{YDGX}^{\ast}$) can be shown by the same argument as in the proof of Theorem 3.2.
@@
where the second and fifth lines follow from Lemma C.1 (i); the third line follows from equation (56); the fourth line follows from the result $F_{Y_{d}SX}=F_{Y_{d}SX}^{\ast}$; the last line follows from equation (54). We thus prove that $\mathbb{E}_{F}[\psi(Y_{1},Y_{0})]=\theta$.
Consequently, we have shown that $F\in\widetilde{\mathcal{F}}^{\ast}$ and $\mathbb{E}_{F}[\psi(Y_{1},Y_{0})]=\theta$. Since this discussion holds for any $\theta\in\widetilde{\Theta}_{IC}^{\dagger}$, we obtain the result $\widetilde{\Theta}_{IC}^{\dagger}\subseteq\Theta_{IC}^{\dagger}$. ∎

### v4

@@ Appendix C Proofs and a Preliminary Lemma
We will show that $F\in\widetilde{\mathcal{F}}^{\ast}$ and that $\mathbb{E}_{F}[\psi(Y_{1},Y_{0})]=\theta$. For the former, we will specifically show that (i) $F$ satisfies Assumptions 2.12.5 and 5.15.2 (with $F^{\ast}$ replaced by $F$) and that (ii) $F_{YDGX}=F_{YDGX}^{\ast}$.
We begin by verifying condition (i). By the same argument as in the proof of Theorem 3.3, it follows that that $F$ satisfies Assumptions 2.12.5, or equivalently $F\in\mathcal{F}^{\ast}$. Applying Lemma C.1(i) to (56), we obtain
$F_{Y_{1}Y_{0}|SX}(\cdot,\cdot|S,X)=C\left(F_{Y_{1}|SX}^{\ast}(\cdot|S,X),F_{Y_{0}|SX}^{\ast}(\cdot|S,X)\middle|S,X\right)\mbox{ a.s}.$
By the definition of $\widetilde{\mathcal{C}}_{X}$, $F_{Y_{1}Y_{0}|SX}$ satisfies conditions (52) and (53). Hence, $F$ also satisfies Assumptions 5.1 and 5.2. In summary, we have shown that $F$ satisfies condition (i).
Condition (ii), namely $F_{YDGX}=F_{YDGX}^{\ast}$, follows by the same argument as in the proof of Theorem 3.3.
@@
where the second and fourth equalities follow from Lemma C.1(i), the third from (56), and the last from (32). The fifth equality follows because (55) together with Lemma C.1(ii) implies $F_{SX}=F_{SX}^{\ast}$. Thus, we have established that $\mathbb{E}_{F}[\psi(Y_{1},Y_{0})]=\theta$.
Consequently, we have shown that $F\in\widetilde{\mathcal{F}}^{\ast}$ and $\mathbb{E}_{F}[\psi(Y_{1},Y_{0})]=\theta$; thus, $\theta\in\Theta_{IC}^{\dagger}$. Since this argument holds for any $\theta\in\widetilde{\Theta}_{IC}^{\dagger}$, we conclude that $\widetilde{\Theta}_{IC}^{\dagger}\subseteq\Theta_{IC}^{\dagger}$. ∎